OpenStudy (the_reginator):

Simplify completely the quantity x squared plus x minus 12 over quantity x squared minus x minus 20 divided by the quantity 3 x squared minus 24 x plus 45 over quantity 12 x squared minus 48 x minus 60.

OpenStudy (danjs):

Is this what it looks like? \[\large \frac{ \frac{ x^2+x-12 }{ x^2-x-20 } }{ \frac{ 3x^2-24x+45 }{ 12x^2-48x-60 } }\]

OpenStudy (the_reginator):

Yes

OpenStudy (danjs):

first you can flip the bottom fraction over and multiply instead of divide.. \[\large \frac{ \frac{ a }{ b } }{ \frac{ c }{ d } }=\frac{ a }{ b }*\frac{ d }{ c }\]

OpenStudy (danjs):

\[\large \frac{ x^2+x-12 }{ x^2-x-20 }*\frac{ 12x^2-48x-60 }{ 3x^2-24x+45 }\]

OpenStudy (the_reginator):

Ok I get it

OpenStudy (danjs):

Then you can pull a factor of 12 from one and a 3 from another... \[\large \frac{ x^2+x-12 }{ x^2-x-20 }*\frac{ 12(x^2-4x-5) }{ 3(x^2-8x+15) }\]

OpenStudy (danjs):

And factor all four of those things.. \[\large \frac{ (x-3)(x+4) }{ (x-5)(x+4) }*\frac{ 12(x-5)(x+1) }{ 3(x-5)(x-3) }\]

OpenStudy (danjs):

the 12/3 reduces to 4/1 and a bunch of those quantities cancel from top and bottom \[\frac{ 4(x+1) }{ x-5 }\]

OpenStudy (danjs):

Thats all you can do ..

OpenStudy (the_reginator):

Ahh ok. I kept on running into walls trying to solve. Thanks :)

OpenStudy (danjs):

welcome